OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (anonymous):

Can anyone on help me?

OpenStudy (anonymous):

@lalaly

OpenStudy (anonymous):

You would use elimination. Choose two of the equations and eliminate one of the three variables. Then take another two equations (one of them will be from before) and cancel out the same variable. Now you have two equations with two variables. From here you could use either elimination or substitution to find one, then both of the variables in those equations. Once you have those, you would plug them back into the original equation to solve for the variable that you initially cancelled out.

OpenStudy (anonymous):

So I could subtract x from the second/third equation to start off right?

OpenStudy (anonymous):

Yes but you would have to make one of the equations entirely negative so that you would have opposite signs and could subtract.

OpenStudy (anonymous):

That way they would cancel out

OpenStudy (anonymous):

So if I multiplied the second equation by -1, and then subtracted with 3rd equation, I would have y-7z=3

OpenStudy (anonymous):

elimination method .....it would be easy to eliminate z first and continue with two variables ......then y=72/29 x=-130/29 z=-15/29